OpenStudy (anonymous):

Please somebody help me!!! I have to use double angle identity to fill in the missing information. cos350degrees= 1-2______________

zepdrix (zepdrix):

\[\Large\bf\sf \cos2 \color{royalblue}{\theta}\quad=\quad 1-2\sin^2\theta\] \[\Large\bf\sf \cos(350)\quad=\quad \cos(2\cdot \color{royalblue}{175})\]

zepdrix (zepdrix):

Understand how to apply the identity after we split the angle like that?

OpenStudy (anonymous):

Really, thank you i had this answer but i doubted myself. I appreciate it

zepdrix (zepdrix):

Oh ok cool :) Also, \(\Large\bf \color{#008353}{\text{Welcome to OpenStudy! :)}}\)

OpenStudy (anonymous):

so would the answer be 1-2sin^2175degrees Thank for the warm welcome

zepdrix (zepdrix):

ya looks good!

OpenStudy (anonymous):

can I ask a little more please because my teacher went too fast for me to catch the concept

zepdrix (zepdrix):

k

OpenStudy (anonymous):

Oh thank you this is tricky tan(x + pi/4)= (1+tanx)/(1-tanx)

zepdrix (zepdrix):

Solve for x?

OpenStudy (anonymous):

I have to prove the one side equals the other

zepdrix (zepdrix):

Oh oh ok :)

zepdrix (zepdrix):

Here is your `Additive Identity for Tangent`\[\Large\bf\sf \tan(\alpha+\beta)\quad=\quad \frac{\tan \alpha+\tan \beta}{1- \tan \alpha \tan \beta}\] \[\Large\bf\sf \tan\left(x+\frac{\pi}{4}\right)\quad=\quad \frac{\tan x+\tan \frac{\pi}{4}}{1- \tan x \tan \frac{\pi}{4}}\]

zepdrix (zepdrix):

Do you remember your unit circle for tangent(pi/4) ? :)

OpenStudy (anonymous):

sqrt2/2

zepdrix (zepdrix):

Nooo silly. That's what sine and cosine give us (each of them).

zepdrix (zepdrix):

Tangent is sine/cosine.

OpenStudy (anonymous):

oh my bad

OpenStudy (anonymous):

we barely praticed tangent im kinda new to it sorry

zepdrix (zepdrix):

tangent = y/x